User name Seonho Choi
Log entry time 15:30:36 on December 9,2007
Entry number 219399
keyword=Wait and see that He level for L dipole come back to normal
Scott has been paged. He suggested to wait until He level comes to normal. Called MCC and they confirmed that there is no problem with CHL. While waiting, continue taking data with right spectrometer.